OpenStudy (anonymous):

Solve the system of equations x + 3y = 1 2x + y = 23 PLEASE HELP

OpenStudy (luigi0210):

any specific method?

OpenStudy (anonymous):

x+3y=1 First solve for x -3y -3y x=-3y+1 Now sustitue that for x x+(-3y+1)+3y=1 x+1=1 cause the y canceled out now -1 the one and x= 0 Now substitute for x 0+3y=1 /3 y=1/3y So your points are (0,1/3y) do you understand :)

OpenStudy (anonymous):

no just need the answer plz its important

OpenStudy (anonymous):

I gave it already lol its x=0 and y=1/3y

OpenStudy (anonymous):

okay thank you soo much

OpenStudy (anonymous):

lol no prob :) glad i could help